OpenStudy (anonymous):

Use the equation below to find the center and radius of the circle. x2 -2x +y2 -2y -23 =0 a. (-1,-1) r=23 b. (1,1) r= 23 c. (1,-1) r=25 d. (1,1) r =5

OpenStudy (anonymous):

complete the square for both x and y

OpenStudy (anonymous):

help me I don't know how to

OpenStudy (anonymous):

for x you have x^2-2x= \[(x ^{2}-2x+1)-1\]

OpenStudy (anonymous):

that is \[(x-1)^{2}-1\] right?

OpenStudy (anonymous):

similarly for y y^2-2y=(y-1)^2 -1

OpenStudy (anonymous):

then \[(x-1)^{2}-1+(y-1)^{2}-1-23=0\]

OpenStudy (anonymous):

can you try to continue ?

OpenStudy (anonymous):

I can try

OpenStudy (anonymous):

I really don't get the problem

OpenStudy (anonymous):

ok collect -1-1-23=-25 and then you ahve (x-1)^2 +(y-1)^2=25 right?

OpenStudy (anonymous):

ok im trying to find the radius and the center though